OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find [f(4)] 2. 81 257 289

OpenStudy (anonymous):

x^2 and the 2 at the end being the same.

OpenStudy (anonymous):

so is it f(x)=x^2+1 and g(x)=3x+1, find [f(4)]^2

OpenStudy (anonymous):

yes

OpenStudy (ranga):

Can you tell me what is f(4) when f(x) = x^2 + 1?

OpenStudy (anonymous):

no? I don't know how to do this.

OpenStudy (anonymous):

take 4 and plug it in the place of x

OpenStudy (ranga):

Just put x = 4 in f(x) and find the value

OpenStudy (anonymous):

this is funtions

OpenStudy (ranga):

f(x) = x^2 + 1 f(4) = 4^2 + 1 = ?

OpenStudy (anonymous):

i don't know..

OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find [f(4)] 2. 81 257 289 Those are my three options.

OpenStudy (ranga):

Looks like you just want the answer. I am sorry.

OpenStudy (anonymous):

I'm on my last attempt of this whole lesson and I am stuck on just this question

OpenStudy (anonymous):

what is 4^2?

OpenStudy (anonymous):

4^2 means 4 squared.

OpenStudy (anonymous):

yea, what does 4 squared equal?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

yes, so f(4)= 4^2+1 f(4)=16+1 =?

OpenStudy (anonymous):

that doesn't equal any of my options though.

OpenStudy (anonymous):

remember you have to square f(4)

OpenStudy (anonymous):

so 33?

OpenStudy (anonymous):

no, what did you get for f(4)=16+1

OpenStudy (anonymous):

33?

OpenStudy (anonymous):

16+1=?

OpenStudy (anonymous):

is 17

OpenStudy (anonymous):

yes so f(4)=17 so take 17 and square it (17)^2=?

OpenStudy (anonymous):

289

OpenStudy (anonymous):

and is that one of your options

OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find 2f(1) + 3g(4) help me with this question now?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

x2 meaning x^2

OpenStudy (anonymous):

for f(1), you would take 1 and plug it in the place of x for the function f(x) = x^2 + 1

OpenStudy (anonymous):

show me how to plug it in

OpenStudy (anonymous):

so f(1)=1^2+1?

OpenStudy (anonymous):

yup now what does f(1)=1^2+1=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

no, what is does 1^2=?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

yea, so f(1)=1^2+1=?

OpenStudy (anonymous):

1+1+1 is 3?

OpenStudy (anonymous):

no, so 1^2= 1 times 1=1 f(1)=1^2+1 f(1)=1+1=?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

what does 1+1=?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

yea so f(1)=1^2+1=?

OpenStudy (anonymous):

I'm so confused. is it 5 or 6?

OpenStudy (anonymous):

f(x)=x^2+1 f(1)=1^2+1 f(1)=1+1 f(1)=?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

now what's the next step from 2?

OpenStudy (anonymous):

so since it is 2f(1) what would 2 times f(1)=? knowing f(1)=2

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so thats half of the problem done 2f(1) + 3g(4) 4 + 3g(4)

OpenStudy (anonymous):

now for g(x)=3x+1 what would you plug in for x, knowing that you are looking for g(4)

OpenStudy (anonymous):

a 4

OpenStudy (anonymous):

yup, what would it look like

OpenStudy (anonymous):

g(4)= 3(4)+1?

OpenStudy (anonymous):

yup, now what does that equal

OpenStudy (anonymous):

13 + 12?

OpenStudy (anonymous):

no, so 3(4)=?

OpenStudy (anonymous):

12/

OpenStudy (anonymous):

12+1=?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

so g(4)= 3(4)+1?

OpenStudy (anonymous):

what would it equal

OpenStudy (anonymous):

13 + 14?

OpenStudy (anonymous):

or 13 + 13?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

g(x)=3x+1 g(4)=3(4)+1 g(4)=12+1 g(4)=?

OpenStudy (anonymous):

g(4)= 13

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so if you look back it says to find 3g(4) 3 times g(4)=? knowing that g(4)=13

OpenStudy (anonymous):

13*3 is 39?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now 2f(1) + 3g(4)=?

OpenStudy (anonymous):

4+39?

OpenStudy (anonymous):

43

OpenStudy (anonymous):

yup WOOOOHOOOOO

OpenStudy (anonymous):

I hate ALG 2, my teacher doesn't know how to teach

OpenStudy (anonymous):

ohhh that sucks

OpenStudy (anonymous):

you can message me if you have questions

OpenStudy (anonymous):

Thanks.